OpenStudy (anonymous):

let h=Vo^2/4.9 sin(theta)cos(theta) model the horizontal distance in meters traveled by a projectile. If the initial velocity is 44 meters/second, which equation would you use to find the angle needed to travel 150 meters? A. 395.10sin(2theta)=150 B. 8.98sin(2theta)=150 C. 150sin(2theta)=150 D, 197.55sin(2theta)=150

OpenStudy (anonymous):

D because 150 =(44)^2/4.9 ( 1/2sin2x) since sin2x=2sinxcosx divide by 2 150=44.44/4.9 .1/2 sin2x 150=22.44/4.9sin2x =968/4.9 sin2x (use calculater) to find it D